The total pressure in a tire is 1.40 atm at 23.0∘C. If the temperature is increased to 32.0∘C, and we assume the tire volume does not change, what is the new pressure?
Answer #1

By Gay Lennacs law, I = 22 25 Here, P, = 1.400m, P2= ? T, = 23°C=(23+473)K = 296K T2 - 32°C = (32+273) K = 305 K By putting